OpenStudy (love_to_love_you):

Which of these is a critical point of the function f(x) = 1/2 x^4 - 3x^3 + 1/2 x^2 + 12x + 7 in the interval [-1.5, 2]?

OpenStudy (jango_in_dtown):

First find the derivative of f and then set it to zero. solve and get the values of x and you have the critical points

OpenStudy (jango_in_dtown):

f'(x)=??

OpenStudy (welshfella):

differentiate term by term using the rule if f(x) = ax^n , f'(x) = anx^(n-1)

OpenStudy (3mar):

Critical points of f(x) are when f'(x)=0 or when f'(x) is undefined

OpenStudy (3mar):

@love_to_love_you Do you follow?

OpenStudy (love_to_love_you):

finally, now I can read y'all's responses. Sorry, my computer wasn't working

OpenStudy (love_to_love_you):

-1.5 -1 1 1.25 Those are the choices, I graphed the function, is there any way I can determine what the critical point is through the graph?

OpenStudy (welshfella):

Yes the critical point will be a turning point on the graph

OpenStudy (solomonzelman):

Suppose (for instance) that you want to find the critical points of \(\color{black}{f(x)=(1/3)x^3-(5/2)x^2+6x-4}\) over the interval \(\color{black}{x \in (e,10]}\). THEORY: Critical points of any function \(\color{black}{f(x)}\) (in general) are possible in 3 cases: \(\color{black}{\small [1]}\) If \(\color{black}{f'(a)=0}\), then \(\color{black}{f'(x)}\) has a critical point (CP) at \(\color{black}{x=a}\). \(\color{black}{\small [2]}\) If \(\color{black}{f'(a)}\) is undefined (i.e. DNE) and \(\color{black}{f(a)}\) exists, then \(\color{black}{f'(x)}\) has a CP at \(\color{black}{x=a}\). \(\color{black}{\small [3]}\) Interval's closed boundary is a CP. SOLUTION: \(\color{black}{f(x)=(1/3)x^3-(5/2)x^2+6x-4}\) \(\color{black}{f'(x)=x^2-5x+6}\) \(\color{black}{0=x^2-5x+6}\) \(\color{black}{0=(x-2)(x-3)}\) So, \(\color{black}{x=2}\) and \(\color{black}{x=3}\). (Note that our interval starts from \(\color{black}{x=e}\) so the only critical point for \(\color{black}{f(x)}\) (thus far) is \(\color{black}{x=3}\). Note also that \(\color{black}{f(x)}\), and by definition (as well as by common sense, and all world's logic) polynomials are always differentiable. (In other words, \(\color{black}{f'(x)}\) is defined (as well as differentiable) everywhere). So, case [2] excluded. Now, we had two interval boundaries, and since \(\color{black}{x=10}\) is the only included boundary (only closed boundary) your second critical point is \(\color{black}{x=10}\). So, the answer is: \(\color{black}{x=3,~10}\).

OpenStudy (solomonzelman):

I meant to say that ... note also that f(x) is a polynomial.

OpenStudy (solomonzelman):

If you have any questions (conceptual, practical or any), please ask and we will address them.

OpenStudy (solomonzelman):

If you haven't learned calculus, then the only way is graphing. Closed interval boundaries, sharp angles (within the interval, if any is given), and points of horizontal tangencies, are all the critical points.

OpenStudy (love_to_love_you):

On my graph, (-1, -1) is a turning point. Would -1 be my answer?

OpenStudy (solomonzelman):

there are two other solutions.

OpenStudy (solomonzelman):

Yes, (-1,-1) is correct.

OpenStudy (solomonzelman):

oh my bad, I didn't look at the interval.

OpenStudy (solomonzelman):

this is the only solution ... good job.

OpenStudy (love_to_love_you):

Thank you

OpenStudy (love_to_love_you):

Thanks everyone for helping

OpenStudy (solomonzelman):

\(y \omega\)